# The SRG/eROSITA All-Sky Survey: X-ray beacons at late cosmic dawn

### v1

@@ 2.1.3 Sample cleaning
We have matched our quasar candidate sample of 2605 sources to compilations of known high-redshift quasars and UCDs. We had blindly selected 27 previously confirmed quasars at $6.85.7<z<$ 6.8 from Mortlock et al. (2009); Jiang et al. (2008); Bañados et al. (2016); Reed et al. (2017); Decarli et al. (2018); Chehade et al. (2018); Reed et al. (2019); Eilers et al. (2020); Venemans et al. (2020); Matsuoka et al. (2022) and Banados et al. (2022), i.e. 77 % of all currently confirmed quasars in the joint DES, VHS and CatWISE2020 surveys. Additionally, our sample contained 7 known M, L and T dwarfs from Liu et al. (2002); Tinney et al. (2018) and Best et al. (2018). These sources are discarded. We also removed one known low-redshift galaxy from Colless et al. (2001) from our sample.
We further extracted forced photometry on optical DES images of the quasar candidates to identify artifacts and problematic blue-band PSF matched photometry (e.g. Bañados et al. 2016; Banados et al. 2022). Circular apertures of radius 3” were centered on the optical coordinates of the candidates. We discard candidates for which we measure a significant aperture flux in g-band, i.e. with errors on g-band aperture magnitudes $\mathtt{ap\_magerr\_g}>0.36$ when a positve g-band flux is measured. Given the size of the source extraction region compared to the size of the DES point-spread-function (PSF $<1.11^{\prime\prime}$, Abbott et al. 2021) this step effectively removes photometric dropouts that lie close to (or in close pairs with) g-band detected stars and lower-redshift galaxies. Example images of discarded candidates are presented in Appendix A. Following this step and a careful visual inspection of each candidate we conserve a sample of 2098 quasar candidates.
@@
We search for X-ray luminous quasars in eRASS (Merloni et al. 2024) and we therefore only account for quasar candidates contained in the intersection region of the optical/IR and eRASS footprints (1604 candidates). The combined area of our survey is measured with a coarse multi-order coverage using the python package pymoc. It spans $3507\,\mathrm{deg^{2}}$. The survey footprint and the on-sky distribution of quasars discovered in this work is shown in Fig. 4.
We verified whether our candidates were X-ray detected in any of the eRASS scans (eRASS1-4) or the deepest, complete cumulative survey eRASS:4. The eROSITA data analyzed in this work were processed with the standard processing pipeline (Brunner et al. 2022, processing version 020). We performed forced photometry on the the eRASS1-4 and eRASS:4 maps at the coordinates of the quasar candidates following a method presented by Arcodia et al. (2024). In each eRASS scan and the deepest available cumulative image (up to eRASS:4) we extracted counts within a circular source aperture of radius 30”, corresponding the the half-energy width (HEW) of the of the eROSITA PSF in the most sensitive energy range (0.2-2.3 keV Dennerl et al. 2012; Predehl et al. 2021; Brunner et al. 2022; Merloni et al. 2024). We note, however, that the HEW of the PSF increases to 34.4” in the 2.3-5.0 keV range. Additionally background photons are extracted in an annulus of standard radii 4 times and 12 times the source region radius. Significantly detected sources in the background region are masked. The binomial no-source probability $P_{B}$ (e.g. Weisskopf et al. 2007; Vito et al. 2019) is given by:
@@
Here $s$ and $b$ are counts within the source and background region, while r is given by the ratio of areas of these two extraction regions. We set the threshold to be less than $0.7\%$ in the wavelength range 0.2-2.3 keV, were eROSITA is most sensitive. According to the eRASS1 simulations of Seppi et al. (2022), this corresponds to a maximal fraction of $\sim 25\%$ of spurious X-ray detections of background fluctuations. We note that this threshold is looser than the one applied by Arcodia et al. (2024, $P_{B}<0.0003$), which is expected to correspond to a spurious fraction of $1\%$. However, assuming typical X-ray to optical flux ratios for Type 1 AGN (e.g. Steffen et al. 2006) $z>5.6$ quasars are expected to lie right at the sensitivity limit of the survey (Wolf et al. 2021) and we therefore apply a more generous threshold. The fraction of spurious background fluctuation detections in eRASS, $f_{\mathrm{bkg}}$, predicted by (Seppi et al. 2022) is shown as a function of $P_{B}$ in Fig. 5.
Due to the large extraction radius of the rigid source region aperture, a caveat of our detection method is the non-negligible possibility of chance alignment, i.e. the astrometric overlap of the extraction region centered on the optical coordinates of a quasar candidate and an unrelated, neighboring eRASS source. The issue of such spurious matches in cross-survey catalogue matching has led to the development of probabilistic approaches (e.g. Sutherland & Saunders 1992; Brusa et al. 2007; Budavári & Szalay 2008; Salvato et al. 2018, 2022). We first estimate the probability of a fully spurious association between a quasar candidate and a single-eRASS (or an eRASS:4) X-ray source, based on the eROSITA catalogues. The eRASS1 source catalogue contains 1,277,477 sources. Here we have also included sources from the supplementary catalogue, with a detection likelihood threshold $L=-\mathrm{ln}(P)\geq 5$ in the 0.2-2.3 keV band (Merloni et al. 2024). Here $P$ is the probability of a source arising from a random background fluctuation (Brunner et al. 2022), i.e. it is directly comparable to the binomial no-source probability $P_{B}$ of Eq. 1. We count 302,161 eRASS1 sources in the common DES, VHS, CatWISE2020 and eRASS1 footprint. Following (Pineau et al. 2017), assuming that the quasar candidates and the eRASS1 sources are completely unrelated samples and using circular positional uncertainties, the total number of fully random associations is given by:
@@
where $\Omega_{\mathrm{e}RASS}(P_{B})=\pi\;\sigma_{\mathrm{eRASS}}(P_{B})^{2}$ and $\Omega_{\mathrm{o}pt}=\pi\;\sigma_{\mathrm{opt}}^{2}$. $\sigma_{\mathrm{eRASS}}$ and $\sigma_{\mathrm{opt}}$ are respectively the fixed source extraction radius in the eRASS map and the quasar candidates. $\Omega$ is the area of the overlap region of the eRASS and DES DR2 footprints. $n_{\rm eRASS}$ and $n_{\rm opt}$ are the number of eRASS1 sources and candidates in this overlap region. The number of eRASS1 sources considered in the surveyed footprint, as well as typical positional uncertainties, are functions of the detection likelihood (DET_LIKE in the eRASS catalogues), or equivalently the binomial no-source probability $P_{B}$. We therefore bin the the full eRASS1 catalogue in DET_LIKE and compute median uncertainties in each bin from the RADEC_ERR=$\sqrt{\sigma_{\texttt{RA}}^{2}+\sigma_{\texttt{DEC}}^{2}}$ values. As RADEC_ERR is the quadratic sum of the 1-dimensional errors in right-ascension and declination, and we assume circular errors for our calculation ($\sigma_{\texttt{RA}}=\sigma_{\texttt{DEC}}$), we consider error radii $\sigma_{\rm eRASS}$=RADEC_ERR$/\sqrt{2}$ for the eRASS1 sources (Pineau et al. 2017). $n_{\mathrm{spurious}}$ can then be calculated for each DET_LIKE bin. The resulting curve of expected spurious associations is shown in Fig. 5. At $P_{B}\sim 10^{-3}$ less than 10 candidates out of 1604 are expected to be randomly overlapping within 30” with an unrelated eRASS1 source.
@@
Figure 5: Upper panel: The cumulative distribution of expected random (unrelated) X-ray detections from forced photometry using a circular aperture of radius 30” around 1604 optical quasar candidate positions with. The black curve is the estimate obtained from Eq. 2. for a single eRASS survey. The black dashed curve shows an empirical estimate of the same number based on a random sample of positions in the surveyed field. The brown curve shows the number of expected contaminants obtained from a random sample in eRASS:4. Blue vertical lines indicate the $P_{B}$ values of the best detection of quasars discovered in this work in any eRASS and in eRASS:4 (if available). Lower panel: Fraction of background sources $f_{\rm bkg}$ expected to be detected in eRASS1 from a simulation by Seppi et al. (2022).
@@
Table 1: Number of detected quasar candidates in each eRASS
@@ 3 Spectroscopic candidate confirmation
Figure 6: Spectra of newly discovered eRASS quasars (left panel) and eRASS:4 (0.2- 8 keV) postage stamps (right panel). The location of the Lyman $\alpha$ break is indicated by a red dashed line. The cutouts are $2^{\prime}\times 2^{\prime}$ in size and are centred on the optical position of the quasars. The spectrum of J202040-621509 shows prominent broad absorption features. The continuum appears strongly absorbed over the entire spectral range bluewards of Lyman $\alpha$. It appears in the eROSITA images as two distinct sources. The morphology J020916-562650 is more compact.
@@
Figure 7: Spectra of three newly discovered quasars, selected during the calibration of the pipeline. These objects are not robustly detected in eROSITA. The red-dashed line indicates the location of the Lyman $\alpha$ break.
@@
From our 75 candidates, we obtained optical spectroscopy for 2 objects so far.They were selected based on observability from the Las Campanas Observatory. In addition, as this pipeline was developed and calibrated, 7 more targets were submitted for spectroscopic follow-up. All of them were selected as significant X-ray sources in eRASS data processed with an earlier version of the eROSITA reduction pipeline. In the latest processing version (020), they are not detected anymore according to the criterion defined in Section 2.2. This is likely due to random fluctuations and to the fact that these objects lie at the sensitivity limit of the survey.
@@
The targets were observed with the LDSS3-C spectrograph mounted on the Magellan-Clay telescope at the Las Campanas Observatory (Chile). The observations took place on July 22 2021 as well as on 2021, October 12 and 26. Long-slit spectra with slit-width $1^{\prime\prime}$ were obtained using the VPH-Red grism, which spans the wavelength range 6000-10500 Å at a dispersion of about 1.16 Å/px. Total exposures varied between 1800 s to 3600 s. A summary of these observations is given in Table 2. Overscan subtractions, flat-field corrections, wavelength, and flux calibrations were performed with IRAF.
@@
Both of the eRASS-detected quasars candidates are confirmed as high-redshift quasars. Their LDSS3-C discovery spectra and eRASS:4 cutouts are presented in Fig. 6. In addition, 3 of the remaining 7 targets observed with LDSS3-C in the development phase of the pipeline, are also confirmed as high-redshift quasars. The spectra of these objects are presented separately in Fig. 7. The quasars are identified by the clear, abrupt, and near-total absorption of their blue continua, attributable to the Lyman $\alpha$ absorption by the increasing density of reionization-era neutral hydrogen clouds. As the unabsorbed continuum and most emission lines of the quasars are shifted out of the sampled spectral range, redshifts can only be measured from the Lyman absorption break. The measurement precision is affected by the low signal-to-noise of some of the spectra. Absorption redwards of Lyman $\alpha$ such as IGM damping wings (Rybicki & Lightman 1979; Miralda-Escudé 1998) can also severely affect Lyman $\alpha$ redshift measurements. We fit a step function with the python package lmfit (StepModel() + ConstantModel()) to the quasar spectra. We estimate the redshift and its error from the break center by bootstrapping from the flux errors. We obtain values between $z\sim 5.6-6.1$. The remaining 4 targets all have spectra corresponding to brown dwarf templates (the coordinate of these sources are presented in Appendix B). We further extrapolate the restframe absolute magnitude at 1450 $\AA$, M1450, from the reddened y-band magnitudes in the DES DR2 catalogues using the spectral index $a_{\nu}=-0.44$ (Vanden Berk et al. 2001). We note that because of the simultaneous DES, VHS, and CatWISE2020 selection and the respective depth of these surveys, the unabsorbed continuum slopes of the selected quasars are biased towards the red values. The choice of the extrapolation from the y-band mitigates the uncertainty due to the expected scatter in the spectral index, as its central wavelength is relatively close to 1450 $\AA$$\times(1+z)$. M1450 measurements are listed in Table 2.
The quasar J202040-621509 was also selected as a quasar candidate by Ighina et al. (2023) who selected this source based on a radio detection with ASKAP (the radio properties of this source are discussed in Section 4.4). We note, however, that our spectroscopic confirmation of this quasar pre-dates this publication.
@@
J202040-621509 305.1702 -62.2526 2021-07-22 2400 $5.718^{+0.001}_{-0.001}$ $19.2345\pm 0.0073$27.32
J020916-562650 32.3205 -56.4473 2021-10-26 2700 $5.606^{+0.001}_{-0.001}$ $20.8983\pm 0.0248$25.76
J050222-341158 75.5930 -34.1995 2021-10-26 2700 $5.793^{+0.001}_{-0.001}$ $21.5331\pm 0.047$25.26
J230341-542730 345.9209 -54.4585 2021-10-12 2400 $5.711^{+0.004}_{-0.002}$ $19.6143\pm 0.0104$27.06
J050411-254958 76.04958 -25.8330 2021-10-12 3600 $6.072^{+0.002}_{-0.002}$ $21.3531\pm 0.0406$25.89
Table 2: Summary of the 5 newly discovered quasars. [1] Coordinates of the centroid of the optical source. [2] Observation date [3] Exposure time of the spectroscopic observation with LDSS3-C [4] Spectroscopic redshift [5] DES z-band magnitude [6] restframe 1450 $\AA$ absolute magnitude.

### v2

@@ 2.1.3 Sample cleaning
We have matched our quasar candidate sample of 2605 sources to compilations of known high-redshift quasars and UCDs. We had blindly selected 27 previously confirmed quasars at $6.85.7<z<$ 6.8 from Mortlock et al. (2009); Jiang et al. (2008); Bañados et al. (2016); Reed et al. (2017); Decarli et al. (2018); Chehade et al. (2018); Reed et al. (2019); Eilers et al. (2020); Venemans et al. (2020); Matsuoka et al. (2022) and Bañados et al. (2023), i.e. 77 % of all currently confirmed quasars in the joint DES, VHS and CatWISE2020 surveys. Additionally, our sample contained 7 known M, L and T dwarfs from Liu et al. (2002); Tinney et al. (2018) and Best et al. (2018). These sources are discarded. We also removed one known low-redshift galaxy from Colless et al. (2001) from our sample.
We further extracted forced photometry on optical DES images of the quasar candidates to identify artefacts and problematic blue-band PSF-matched photometry (e.g. Bañados et al. 2016, 2023). Circular apertures of radius 3” were centred on the optical coordinates of the candidates. We discard candidates for which we measure a significant aperture flux in g-band, i.e. with errors on g-band aperture magnitudes $\mathtt{ap\_magerr\_g}>0.36$ when a positive g-band flux is measured. Given the size of the source extraction region compared to the size of the DES point-spread-function (PSF $<1.11^{\prime\prime}$, Abbott et al. 2021), this step effectively removes photometric dropouts that lie close to (or in close pairs with) the g-band detected stars and lower-redshift galaxies. Example images of discarded candidates are presented in Appendix A. Following this step and a careful visual inspection of each candidate, we conserve a sample of 2098 quasar candidates.
@@
We search for X-ray luminous quasars in eRASS (Merloni et al. 2024), and we therefore only account for quasar candidates contained in the intersection region of the optical/IR and eRASS footprints (1604 candidates). The combined area of our survey is measured with a coarse multi-order coverage using the python package pymoc. It spans $3507\,\mathrm{deg^{2}}$. The survey footprint and the on-sky distribution of quasars discovered in this work are shown in Fig. 4.
We verified whether our candidates were X-ray detected in any of the eRASS scans (eRASS1-4) or the deepest, complete cumulative survey eRASS:4. The eROSITA data analyzed in this work were processed with the standard processing pipeline (Brunner et al. 2022, processing version 020). We performed forced photometry on the eRASS1-4 and eRASS:4 maps at the coordinates of the quasar candidates following a method presented by Arcodia et al. (2024). In each eRASS scan and the deepest available cumulative image (up to eRASS:4), we extracted counts within a circular source aperture of radius 30”, corresponding the half-energy width (HEW) of the eROSITA PSF in the most sensitive energy range (0.2-2.3 keV Dennerl et al. 2012; Predehl et al. 2021; Brunner et al. 2022; Merloni et al. 2024). We note, however, that the HEW of the PSF increases to 34.4” in the 2.3-5.0 keV range. Additionally, background photons are extracted in an annulus of standard radii 4 times and 12 times the source region radius. Significantly detected sources in the background region are masked. The binomial no-source probability $P_{b}$ (e.g. Weisskopf et al. 2007; Vito et al. 2019) is given by:
@@
Here $s$ and $b$ are counts within the source and background region, while r is given by the ratio of areas of these two extraction regions. We set the threshold of $P_{b}$ to be less than $0.7\%$ in the wavelength range 0.2-2.3 keV, where eROSITA is most sensitive. According to the eRASS1 simulations of Seppi et al. (2022), this corresponds to a maximal fraction of $\sim 25\%$ of spurious X-ray detections of background fluctuations. We note that this threshold is looser than the one applied by Arcodia et al. (2024, $P_{b}<0.0003$), which is expected to correspond to a spurious fraction of $1\%$. However, assuming typical X-ray to optical flux ratios for Type 1 AGN (e.g. Steffen et al. 2006) $z>5.6$ quasars are expected to lie right at the sensitivity limit of the survey (Wolf et al. 2021) and we therefore apply a more generous threshold. The fraction of spurious background fluctuation detections in eRASS, $f_{\mathrm{bkg}}$, predicted by (Seppi et al. 2022) is shown as a function of $P_{b}$ in Fig. 5.
Due to the large extraction radius of the rigid source region aperture, a caveat of our detection method is the non-negligible possibility of chance alignment, i.e. the astrometric overlap of the extraction region centred on the optical coordinates of a quasar candidate and an unrelated, neighbouring eRASS source. The issue of such spurious matches in cross-survey catalogue matching has led to the development of probabilistic approaches (e.g. Sutherland & Saunders 1992; Brusa et al. 2007; Budavári & Szalay 2008; Salvato et al. 2018, 2022). We first estimate the probability of a fully spurious association between a quasar candidate and a single-eRASS (or an eRASS:4) X-ray source based on the eROSITA catalogues. The eRASS1 source catalogue contains 1,277,477 sources. Here, we have also included sources from the supplementary catalogue, with a detection likelihood threshold $L=-\mathrm{ln}(P)\geq 5$ in the 0.2-2.3 keV band (Merloni et al. 2024). Here $P$ is the probability of a source arising from a random background fluctuation (Brunner et al. 2022), i.e. it is directly comparable to the binomial no-source probability $P_{b}$ of Eq. 1. We count 302,161 eRASS1 sources in the common DES, VHS, CatWISE2020 and eRASS1 footprint. Following (Pineau et al. 2017), assuming that the quasar candidates and the eRASS1 sources are completely unrelated samples and using circular positional uncertainties, the total number of fully random associations is given by:
@@
where $\Omega_{\mathrm{e}RASS}(P_{b})=\pi\;\sigma_{\mathrm{eRASS}}(P_{b})^{2}$ and $\Omega_{\mathrm{o}pt}=\pi\;\sigma_{\mathrm{opt}}^{2}$. $\sigma_{\mathrm{eRASS}}$ and $\sigma_{\mathrm{opt}}$ are, respectively, the fixed source extraction radius in the eRASS map and the quasar candidates. $\Omega$ is the area of the overlap region of the eRASS and DES DR2 footprints. $n_{\rm eRASS}$ and $n_{\rm opt}$ are the number of eRASS1 sources and candidates in this overlap region. The number of eRASS1 sources considered in the surveyed footprint, as well as typical positional uncertainties, are functions of the detection likelihood (DET_LIKE in the eRASS catalogues), or equivalently the binomial no-source probability $P_{b}$. We therefore bin the full eRASS1 catalogue in DET_LIKE and compute median uncertainties in each bin from the RADEC_ERR=$\sqrt{\sigma_{\texttt{RA}}^{2}+\sigma_{\texttt{DEC}}^{2}}$ values. As RADEC_ERR is the quadratic sum of the 1-dimensional errors in right-ascension and declination, and we assume circular errors for our calculation ($\sigma_{\texttt{RA}}=\sigma_{\texttt{DEC}}$), we consider error radii $\sigma_{\rm eRASS}$=RADEC_ERR$/\sqrt{2}$ for the eRASS1 sources (Pineau et al. 2017). $n_{\mathrm{spurious}}$ can then be calculated for each DET_LIKE bin. The resulting curve of expected spurious associations is shown in Fig. 5. At $P_{b}\sim 10^{-3}$, less than ten candidates out of 1604 are expected to be randomly overlapping within 30” with an unrelated eRASS1 source.
@@
Figure 5: Upper panel: The cumulative distribution of expected random (unrelated) X-ray detections from forced photometry using a circular aperture of radius 30” around 1604 optical quasar candidate positions. The black curve is the estimate obtained from Eq. 2. for a single eRASS survey. The black dashed curve shows an empirical estimate of the same number based on a random sample of positions in the surveyed field. The brown curve shows the number of expected contaminants obtained from a random sample in eRASS:4. Blue vertical lines indicate the $P_{b}$ values of the best detection of quasars discovered in this work in any eRASS and in eRASS:4 (if available). Lower panel: Fraction of background sources $f_{\rm bkg}$ expected to be detected in eRASS1 from a simulation by Seppi et al. (2022).
Table 1: Number of detected quasar candidates in each eRASS
@@ 3 Spectroscopic candidate confirmation
Figure 6: Spectra of newly discovered eRASS quasars (left panel) and eRASS:4 (0.2- 8 keV) postage stamps (right panel). The cutouts are $2^{\prime}\times 2^{\prime}$ in size and are centred on the optical position of the quasars. The spectrum of J202040-621509 shows prominent broad absorption features. The continuum appears strongly absorbed over the entire spectral range bluewards of Lyman $\alpha$. It appears in the eROSITA images as two distinct sources. This morphology is further investigated in Section 4.3.1 with a Chandra follow-up pointing. There we argue that the eRASS source is related to the quasar (Fig 11). The morphology J020916-562650 is more compact.
@@
Figure 7: Spectra of three newly discovered quasars selected during the calibration of the pipeline. These objects are not robustly detected in eROSITA. The red dashed line indicates the location of the Lyman $\alpha$ break.
@@
From our 75 candidates, we obtained optical spectroscopy for two objects so far. They were selected based on observability from the Las Campanas Observatory. In addition, as this pipeline was developed and calibrated, seven more targets were submitted for spectroscopic follow-up. All of them were selected as significant X-ray sources in eRASS data processed with an earlier version of the eROSITA reduction pipeline. In the latest processing version (020), they are not detected anymore according to the criterion defined in Section 2.2. This is likely due to random fluctuations and to the fact that these objects lie at the sensitivity limit of the survey.
@@
The targets were observed with the LDSS3-C spectrograph mounted on the Magellan-Clay telescope at the Las Campanas Observatory (Chile). The observations took place on July 22 2021, as well as on 2021 October 12 and 26. Long-slit spectra with slit-width $1^{\prime\prime}$ were obtained using the VPH-Red grism, which spans the wavelength range 6000-10500 Å at a dispersion of about 1.16 Å/px. Total exposures varied between 1800 s to 3600 s. A summary of these observations is given in Table 4. Overscan subtractions, flat-field corrections, wavelength, and flux calibrations were performed with IRAF.
@@
Both of the eRASS-detected quasars candidates are confirmed as high-redshift quasars [^3] We note that shortly after the submission of the present paper, the independent discovery of quasar J020916-562650 was reported by Ighina et al. (2024). The quasar J202040-621509 was also presented as quasar candidate by Ighina et al. (2023). However, our spectroscopic confirmation of this quasar also pre-dates this publication.. Their LDSS3-C discovery spectra and eRASS:4 cutouts are presented in Fig. 6. In addition, 3 of the remaining 7 targets observed with LDSS3-C in the development phase of the pipeline are also confirmed as high-redshift quasars. The spectra of these objects are presented separately in Fig. 7. The quasars are identified by the clear, abrupt, and near-total absorption of their blue continua, attributable to the Lyman $\alpha$ absorption by the increasing density of reionization-era neutral hydrogen clouds. As the unabsorbed continuum and most emission lines of the quasars are shifted out of the sampled spectral range, redshifts can only be measured from the Lyman absorption break. The measurement precision is affected by the low signal-to-noise of some of the spectra. Absorption effects redwards of Lyman $\alpha$ such as IGM damping wings (Rybicki & Lightman 1979; Miralda-Escudé 1998) can also severely affect Lyman $\alpha$ redshift measurements. We fit a step function with the python package lmfit (StepModel() + ConstantModel()) to the quasar spectra. We estimate the redshift and its error from the break centre by bootstrapping from the flux errors. We obtain values between $z\sim 5.6-6.1$. The remaining 4 targets all have spectra corresponding to brown dwarf templates (the coordinates of these sources are presented in Appendix B). We further extrapolate the restframe absolute magnitude at 1450 $\AA$, M1450, from the reddened y-band magnitudes in the DES DR2 catalogues using the spectral index $a_{\nu}=-0.44$ (Vanden Berk et al. 2001). We note that because of the simultaneous DES, VHS, and CatWISE2020 selection and the respective depth of these surveys, the unabsorbed continuum slopes of the selected quasars are biased towards the red values. The choice of the extrapolation from the y-band mitigates the uncertainty due to the expected scatter in the spectral index, as its central wavelength is relatively close to 1450 $\AA$$\times(1+z)$. M1450 measurements are listed in Table 4.
Table 2: Summary of the 5 newly discovered quasars.
@@
J202040-621509 305.170194 -62.252553 2021-07-22 2400 $5.718^{+0.001}_{-0.001}$ $19.2345\pm 0.0073$27.32
J020916-562650 32.320529 -56.447344 2021-10-26 2700 $5.606^{+0.001}_{-0.001}$ $20.8983\pm 0.0248$25.76
J050222-341158 75.592959 -34.199513 2021-10-26 2700 $5.793^{+0.001}_{-0.001}$ $21.5331\pm 0.047$25.26
J230341-542730 345.920933 -54.458538 2021-10-12 2400 $5.711^{+0.004}_{-0.002}$ $19.6143\pm 0.0104$27.06
J050411-254958 76.049582 -25.832959 2021-10-12 3600 $6.072^{+0.002}_{-0.002}$ $21.3531\pm 0.0406$25.89
444[1] Coordinates of the centroid of the optical source. [2] Observation date [3] Exposure time of the spectroscopic observation with LDSS3-C [4] Spectroscopic redshift [5] DES z-band magnitude [6] restframe 1450 $\AA$ absolute magnitude.